Grade 12Physical Chemistry

a lineman of mass 60 kg is climbing a vertical pole. the co-efficient of static friction between his body and the pole is 0.5.if he is able to climb up the pole,what is minimum force with which he should press the pole?

Nishtha Gahlot
9 Years agoGrade 12

2 Answers

Vikas TU
ApprovedApproved Tutor Answer9 Years ago
Friction force = 0.5*600 => 300N
And Mg force =>  600N
Asthe force is static the net force required would be =>  600 -300 => 300 N.
Sujit Kumar
8 Years ago
I feel the answer by @Vikas TU is incorrect...
My solution:
 
On a vertical pole :
Frictional Force = Weight
Weight = mg = 60 * 10 = 600 N
Frictional Force =  u * Normal Force
 
Therefore, 600 = 0.5 * Normal Force
Which gives, Normal force = 1200N
Minimum force with which he should press the pole = Normal Force = 1200 N